OpenStudy (anonymous):

Melissa has 6 times as many quarters as Michelle. Together, they have a total of 896 quarters. How many quarters, q , does Michelle have?

OpenStudy (anonymous):

x = michelle's quarters m = melissa's quarters m = 6x m + x = 896 --> m = 896 - x now sub 896 - x in for m in the first equation... m = 6x 896 - x = 6x can you solve it from here ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

let me finish then.... 896 - x = 6x (add x to both sides) 896 - x + x = 6x + x 896 = 7x (now divide both sides by 7) 896/7 =7x/7 128 = x now sub 128 in for x in first equation... m = 6x m = 6(128) m = 768 check... m + x = 896 768 + 128 = 896 896 = 896 (correct) Michelle (x) has 128 quarters Melissa (m) has 768 quarters Do you understand how I did this ?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

your welcome :)